OpenStudy (anonymous):

What is the volume of a cone with the given dimensions? Radius = 4 cm ; Height = 10 cm Round the answer to the nearest hundredth. Use 3.14 for pi. a. 167.47 cu. cm b. 112.57 cu. cm c. 484.15 cu. cm d. 964.15 cu. cm

OpenStudy (mathstudent55):

\(\Large V_{cone} = \dfrac{1}{3} \pi r^2 h\) You were given radius = r = 4 cm, height = h = 10 cm pi = \(\pi\) = 3.14 Use those numbers in the formula and find the volume. (Hint: remember to square the radius, as the formula shows.)

OpenStudy (anonymous):

i got 150.72